# Other Vehicle Trajectories Are Also Needed: A Driving World Model Unifies Ego-Other Vehicle Trajectories in Video Latent Space

### v2

@@ Abstract
Advanced end-to-end autonomous driving systems predict other vehicles’ motions and plan ego vehicle’s trajectory. The world model that can foresee the outcome of the trajectory has been used to evaluate the end-to-end autonomous driving system. However, existing world models predominantly emphasize the trajectory of the ego vehicle and leave other vehicles uncontrollable. This limitation hinders their ability to realistically simulate the interaction between the ego vehicle and the driving scenario. In addition, it remains a challenge to match multiple trajectories with each vehicle in the video to control the video generation. To address above issues, a driving World Model named EOT-WM is proposed in this paper, unifying Ego-Other vehicle Trajectories in videos. Specifically, we first project ego and other vehicle trajectories in the BEV space into the image coordinate to match each trajectory with its corresponding vehicle in the video. Then, trajectory videos are encoded by the Spatial-Temporal Variational Auto Encoder to align with driving video latents spatially and temporally in the unified visual space. A trajectory-injected diffusion Transformer is further designed to denoise the noisy video latents for video generation with the guidance of ego-other vehicle trajectories. In addition, we propose a metric based on control latent similarity to evaluate the controllability of trajectories. Extensive experiments are conducted on the nuScenes dataset, and the proposed model outperforms the state-of-the-art method by 30% in FID and 55% in FVD. The model can also predict unseen driving scenes with self-produced trajectories.
@@ 1 Introduction
End-to-end autonomous driving [18, 31, 24] has gained increasing attention recently, since the approach integrates all modules (e.g. perception, decision, and planning modules) into a model optimized jointly to directly output planning results based on input multi-sensor data. Despite the promising performance of the end-to-end autonomous driving model, it remains a challenge to handle out-of-distribution cases, which are probably dangerous and high-cost to simulate as well. The world model [36, 26, 37, 40, 8] can predict the future states of the driving scene based on historical observations and future driving actions, which is a potential solution to evaluate the autonomous driving model and avoid catastrophic errors.
Some world models [8, 40] explore predicting future states of the driving scene with the trajectory, since it can reflect the driving action of the vehicle in the scene more precisely. Despite advanced end-to-end autonomous driving systems (e.g. VAD [19]) can predict other vehicles’ motions and plan ego vehicle’s trajectory, these world models primarily focus on the ego vehicle trajectory. They view the trajectory as a series of points in the bird’s eye view (BEV) space [17, 23], which are directly encoded as the condition for video generation. However, there are three main drawbacks in the approaches mentioned above. Firstly, they just consider the ego vehicle trajectory and leave other vehicles uncontrollable in the generated video. As a result, the model cannot realistically simulate the interaction between the ego vehicle and the driving scenario, and generate diverse novel scenes by changing other vehicle trajectories as well. Secondly, the distribution of the encoding based on the trajectory points in the BEV space is quite different from that of the video latents in visual modality without aligning their feature space. In addition, it is impractical to correspond multiple trajectories in the BEV space to numerous vehicles in visual modality, since the BEV space is mismatched with the video space.
@@ 2.2 World Model
The world model can predict the future states based on historical observations and future actions, which has been extensively applied in simulated games [11, 12, 13] and indoor embodiment [20, 34, 27]. Recently, the world model has attracted great attention in autonomous driving [36, 26, 37, 40, 8]. WoVoGen [26], GenAD [40] and Drive-WM [37] dedicated to generation 6-view driving videos based on the given conditions. Drivedreamer [36] and Vista [8] developed world models to predict future vehicle actions in the front-view video. However, world models mentioned above all concentrated on the actions of the ego vehicle and ignored those of the other vehicles, making the interaction with the environment insufficient in these world models. In this paper, the proposed EOT-WM use the ego and other vehicle trajectories to represent the motions of vehicles in the scene, and generate more realistic videos with all vehicles controllable.
@@ 3.3 Trajectory-injected Diffusion Transformer
Therefore, we design Trajectory-injected Diffusion Transformer (TiDiT) to predict future driving scenes based on specific initial frames and trajectories. Specifically, to enable the model to generate the video from historical $T_{c}$ frames, we replace the first $T_{c}$ frames in noisy video latents $z_{vid}^{noise}$ with the first $T_{c}$ frames of $z_{vid}$. The process can be realize via a conditioning mask $M=\{t\leq T_{c}|1\leq t\leq T\}$, which is formulated as
@@
The original Diffusion Transformer in CogvideoX [41] use the text as the condition to control video generation. To fully reserve the ability of CogvideoX, we utilize the scene caption provided in OmniDrive [35] to train the proposed TiDiT. As a result, visual latents $z_{vid}^{f}$ and text embedding $z_{text}$ extracted from the scene caption are concatenated and fed into several layers of Expert Diffusion Transformer in CogvideoX to predict the noise added to $z_{vid}$ as follow:
@@
In the inference stage, we use the STVAE to extract latents of the initial frames, and replace the corresponding frames of the random sampling noise to obtain $z$. Then, following Eq. 6 and Eq. 7, the Expert Diffusion Transformer inferences the noise for the VAE decoder to generate video.
@@
Then, we adopt two-layer MLP to decode the predicted video latents as the predicted trajectory video latents $z_{traj}^{p}$ as follow
@@
which also serve as the metric to evaluate trajectory controllability. Moreover, the predicted trajectory video latents $z_{traj}^{p}$ can be decoded via STVAE to generate trajectory video for visualization.
Methods Resolution FID $\downarrow$ FVD $\downarrow$ DriveGAN (2016) [30] 256×256 73.4 502.3 DriveDreamer (2023) [36] 128×192 52.6 452.0 WoVoGen (2024) [26] 256×448 27.6 417.7 Drive-WM (2024) [37] 192×384 15.8 122.7 GenAD (2024) [40] 256×448 15.4 184.0 Vista (2024) [8] 576×1024 6.9 89.4 EOT-WM 768×1280 4.8 40.0
@@ 4.1 Dataset and Evaluation Metric
Extensive experiments are conducted on the nuScenes dataset [4] to evaluate the proposed EOT-WM. Following the setup of Vista [8], the training and evaluation set contain 25109 and 5369 videos with 25 frames. The ego vehicle trajectory corresponding to the video can be directly obtained in the annotations. The other vehicle trajectory is only annotated in key frames, and we use interpolation to complete the trajectory corresponding to the video. We use FID [14] and FVD [32] to evaluate the quality of generated videos. Moreover, we provide qualitative demonstrations that underscore the advancements of the proposed EOT-WM.
@@ 4.2 Implementation Details
We bulid the proposed EOT-WM based on CogvideoX-2B [41]. We use the first video latent as the context for predicting the future frames. The frame rate of the generated video is 10Hz. We use the AdamW [25] optimizer with a learning rate of $2\times 10^{-5}$ to train the model. By default, the proposed model is trained at 768×1280 resolution for 60 epochs with 64 NVIDIA A800 GPUs with a total batch size of 128.
@@ 4.3 Comparison with State-of-the-art Methods
Quantitative Result. Comparison results of FID and FVD for video generation achieved by the proposed EOT-WM and competing methods on the validation set of nuScenes dataset are given in Tab. 1. As can be observed, the state-of-the-art driving world model Vista [8] can generate videos with higher quality and resolution compared with previous works [30, 36, 26, 37, 40]. However, Vista [8] cannot control other vehicles in the scene, and using trajectories in BEV space is not sufficiently effective. Therefore, the proposed EOT-WM controlling video generation with ego and other vehicle trajectories in video space achieves 4.8 FID and 40.0 FVD, surpassing Vista by 30% in FID and 55% in FVD. Moreover, the proposed EOT-WM can generate videos in 768×1280 resolution, which is superior to previous world models as well.
Qualitative Result. To further demonstrate the superiority in action controllability of the proposed EOT-WM, we provide several representative cases achieved by the proposed EOT-WM compared with Vista [8] in Fig. 4. As is shown in Fig. 4 (a), the blank car in the red rectangle stands by the side of the road. Vista and EOT-WM w/o O.V. Traj. do not consider the other vehicle trajectories, and the blank car moves forward in the generated videos. In contrast, the blank car keeps still in the video generated by the proposed EOT-WM. In Fig. 4 (b), the truck is backing up on the road. However, Vista and EOT-WM w/o O.V. Traj. both generate the truck moving forward in the generated video. As for the proposed EOT-WM, the action of the truck is predicted correctly in the generated video. These cases demonstrate that the proposed EOT-WM can
@@
Table 3: Effect of difference texts used in the proposed EOT-WM.
generate more realistic videos with ego and other vehicle trajectories.
4.4 Ablation Study
To verify the effectiveness of the proposed modules in EOT-WM, ablation study for key components is conducted on the validation set of nuScenes dataset, and the results are given in Tab. 2. When all components are not used, the model is close to action-free only with initial frames as the context. As a result, the model obtains poor performance as 61.4 FID and 372.8 FVD. As for EOT-WM w/o VTR model, the performance is still poor, since the model cannot learn trajectories in BEV space well and other modules designed for VTR cannot work. For EOT-WM w/o AMGG model, we use C3D to extract trajectory video features, and the performance declines compared with the proposed EOT-WM due to the dissimilarity of feature space. The ego vehicle trajectory is the most important condition for video generation. Therefore, EOT-WM w/o E. V. Traj. model cannot generate realistic videos. With other vehicle trajectories to control the other vehicles in the scene, the proposed EOT-WM model can achieve better FID and FVD compared with EOT-WM w/o O. V. Traj. model. All above results demonstrate the effectiveness of the proposed modules.
4.5 Effect of difference text types
The original CogvideoX [41] is trained to use the text to control video generation, and other advanced video generation models [10, 21] are also text-to-video models. When building the world model based on such text-to-video models, it is unadvisable to remove the text element since the capability of the original text-to-video model will be damaged. As can be observed in Tab. 3, while the proposed EOT-WM does not use the text, FID and FVD achieved decline obviously. Then, we adopt three types of texts to control the video generation model. The performance obtained by these types of texts are similar, demonstrating that the video generation model can be trained to adapt different types of texts. Among them, EOT-WM using the action description achieves the bEOT performance since the caption contains the action information. However, considering the capability of the model to generate novel scenes, we choose the scene description for EOT-WM as the action description will conflict with the novel trajectories.
4.6 Generating novel scenes
To verify the capability of the proposed EOT-WM to generate novel scenes, several instances generated with novel trajectories are given in Fig. 6. As can be observed in Fig. 6 (a), the ego vehicle turns right and the bus moves forward in the video generated with groundtruth ego and other vehicle trajectories in the dataset. While we use self-produced ego vehicle trajectory that represents turning left in Fig. 6 (b), the ego vehicle is controlled to turn left in the video generated. Moreover, we give novel other vehicle trajectory meaning stop in Fig. 6 (c), the bus in the generated video stands by. Above results demonstrate that the proposed EOT-WM can generating novel scenes according to novel trajectories.
4.7 Evaluating Trajectory Controllability
To evaluate trajectory controllability, we propose a metric based on control latent similarity in the video latent space. As is shown in Fig. 7, we use the conditional trajectory video to generate predicted trajectory video latents in TiDiT. The predicted trajectory video latents can be decode into predicted trajectory video, which is similar as the conditional trajectory video. Moreover, given another irrelevant trajectory video, the metric calculated between the predicted trajectory video latents and the conditional trajectory video latents is obviously smaller than that calculated between the predicted trajectory video latents and the irrelevant trajectory video latents. These results demonstrate the effectiveness of the proposed latent-based metric.
@@ 5 Conclusion
In the future work, we will explore the manner to control the other vehicles more precisely and the quantitative metric to evaluation the consistency between the given trajectory and the generated video.

### v3

@@ Abstract
Advanced end-to-end autonomous driving systems predict other vehicles’ motions and plan ego vehicle’s trajectory. The world model that can foresee the outcome of the trajectory has been used to evaluate the autonomous driving system. However, existing world models predominantly emphasize the trajectory of the ego vehicle and leave other vehicles uncontrollable. This limitation hinders their ability to realistically simulate the interaction between the ego vehicle and the driving scenario. In this paper, we propose a driving World Model named EOT-WM, unifying Ego-Other vehicle Trajectories in videos for driving simulation. Specifically, it remains a challenge to match multiple trajectories in the BEV space with each vehicle in the video to control the video generation. We first project ego-other vehicle trajectories in the BEV space into the image coordinate for vehicle-trajectory match via pixel positions. Then, trajectory videos are encoded by the Spatial-Temporal Variational Auto Encoder to align with driving video latents spatially and temporally in the unified visual space. A trajectory-injected diffusion Transformer is further designed to denoise the noisy video latents for video generation with the guidance of ego-other vehicle trajectories. In addition, we propose a metric based on control latent similarity to evaluate the controllability of trajectories. Extensive experiments are conducted on the nuScenes dataset, and the proposed model outperforms the state-of-the-art method by 30% in FID and 55% in FVD. The model can also predict unseen driving scenes with self-produced trajectories.
@@ 1 Introduction
End-to-end autonomous driving [20, 33, 26] has gained increasing attention recently, since the approach integrates all modules (e.g. perception, decision, and planning modules) into a model optimized jointly to directly output planning results based on input multi-sensor data. Despite the promising performance of the end-to-end autonomous driving model, effectively handling out-of-distribution scenarios continues to be a significant challenge, particularly since such situations are often hazardous and costly to simulate. The world model [38, 39, 9, 15] can predict future driving scenes based on historical observations and future driving actions, which is a potential solution to evaluate the autonomous driving model and avoid catastrophic errors.
Some world models [9, 42] explore predicting future driving scenes with the trajectory, since it can reflect the driving action of the vehicle in the scene more precisely. Despite advanced end-to-end autonomous driving systems (e.g. VAD [21]) can predict other vehicles’ motions and plan ego vehicle’s trajectory, these world models primarily focus on the ego vehicle trajectory. They view the trajectory as a series of points in the bird’s eye view (BEV) space [19, 25], which are directly encoded as the condition for video generation. However, there are three main drawbacks in the approaches mentioned above. Firstly, they just consider the ego vehicle trajectory and leave other vehicles uncontrollable in the generated video. As a result, the model cannot realistically simulate the interaction between the ego vehicle and the driving scenario, and generate diverse novel scenes by changing other vehicle trajectories as well. Secondly, the distribution of the encoding based on the trajectory points in the BEV space is quite different from that of the video latents in visual modality without aligning their feature space. In addition, it is impractical to correspond multiple trajectories in the BEV space to numerous vehicles in visual modality, since the BEV space is mismatched with the video space.
@@ 2.2 World Model
The world model can predict the future states based on historical observations and future actions, which has been extensively applied in simulated games [12, 13, 14] and indoor embodiment [22, 36, 29]. Recently, the world model has attracted great attention in autonomous driving [38, 28, 39, 42, 9, 15, 1]. WoVoGen [28] and Drive-WM [39] are devoted to generating 6-view driving videos based on the given conditions. Drivedreamer [38], GenAD [42], Vista [9], and GEM [15] develop world models to predict future driving scenes via the front-view video. However, most world models mentioned above concentrate on the actions of the ego vehicle and ignore those of the other vehicles, making the interaction with the environment insufficient in these world models. Although GEM [15] uses future object features and human poses to generate future videos, it is more impractical for autonomous driving systems to obtain them compared with future trajectories. Particularly, Cosmos [1] is a commercial model comprising over 7 billion parameters, which imposes substantial computational overhead during both training and inference. In this paper, the proposed EOT-WM uses the ego-other vehicle trajectories to represent the motions of vehicles in the scene, and generate more realistic videos with all vehicles controllable.
@@ 3.3 Trajectory-injected Diffusion Transformer
Therefore, we design Trajectory-injected Diffusion Transformer (TiDiT) to predict future driving scenes based on specific initial frames and trajectories. Specifically, to enable the model to generate the video from historical $T_{c}$ frames, we replace the first $T_{c}$ frames in noisy video latents $z_{vid}^{noise}$ with the first $T_{c}$ frames of $z_{vid}$. The process can be realized via a conditioning mask $M=\{t\leq T_{c}|1\leq t\leq T\}$, which is formulated as
@@
The original Diffusion Transformer in CogvideoX [43] uses the text as the condition to control video generation. To fully reserve the ability of CogvideoX, we utilize the scene caption provided in OmniDrive [37] to train the proposed TiDiT. As a result, visual latents $z_{vid}^{f}$ and text embedding $z_{text}$ extracted from the scene caption are concatenated and fed into several layers of Expert Diffusion Transformer in CogvideoX to predict the noise added to $z_{vid}$ as follows:
@@
In the inference stage, we use the STVAE to extract latents of the initial frames, and replace the corresponding frames of the random sampling noise to obtain $z$. Then, following Eq. 6 and Eq. 7, the Expert Diffusion Transformer infers the noise for the VAE decoder to generate video.
@@
Then, we adopt a two-layer MLP to decode the predicted video latents as the predicted trajectory video latents $z_{traj}^{p}$ as follows
@@
which also serves as the metric to evaluate trajectory controllability. Moreover, the predicted trajectory video latents $z_{traj}^{p}$ can be decoded via STVAE to generate trajectory video for visualization.
@@ 4.1 Dataset and Evaluation Metric
Extensive experiments are conducted on the nuScenes dataset [5] to evaluate the proposed EOT-WM. Following the setup of Vista [9] in the nuScenes dataset, the training and evaluation set contain 25109 and 5369 videos with 25 frames. The ego vehicle trajectory corresponding to the video can be directly obtained in the annotations. The other vehicle trajectory is only annotated in key frames, and we use interpolation based on the annotated trajectories to complete the full trajectory corresponding to the 25-frame video. We use FID [16] and FVD [34] to evaluate the quality of generated videos. Moreover, we provide qualitative demonstrations that underscore the advancements of the proposed EOT-WM.
@@ 4.2 Implementation Details
We build the proposed EOT-WM based on CogvideoX-2B [43]. We use the first video latent as the context for predicting the future frames. The frame rate of the generated video is 10Hz. We use the AdamW [27] optimizer with a learning rate of $2\times 10^{-5}$ to train the model. By default, the proposed model is trained at 768×1280 resolution for 60 epochs with 64 NVIDIA A800 GPUs with a total batch size of 128.
@@ 4.3 Comparison with State-of-the-art Methods
Quantitative Result. Comparison results of FID and FVD for video generation achieved by the proposed EOT-WM and competing methods on the validation set of nuScenes dataset are given in Tab. 1. As can be observed, the state-of-the-art driving world model Vista [9] can generate videos with higher quality and resolution compared with previous works [32, 38, 28, 39, 42]. However, Vista [9] cannot control other vehicles in the scene, and using trajectories in BEV space is not sufficiently effective. Therefore, the proposed EOT-WM controlling video generation with ego and other vehicle trajectories in video space achieves 4.8 FID and 40.0 FVD, surpassing Vista by 30% in FID and 55% in FVD. Moreover, the proposed EOT-WM can generate videos in 768×1280 resolution, superior to previous world models as well.
@@
Methods Resolution FID $\downarrow$ FVD $\downarrow$ DriveGAN [32] 256×256 73.4 502.3 DriveDreamer [38] 128×192 52.6 452.0 WoVoGen [28] 256×448 27.6 417.7 Drive-WM [39] 192×384 15.8 122.7 GenAD [42] 256×448 15.4 184.0 Vista [9] 576×1024 6.9 89.4 EOT-WM 768×1280 4.8 40.0
Qualitative Result. To further demonstrate the superiority in action controllability of the proposed EOT-WM, we provide several representative cases achieved by the proposed EOT-WM compared with Vista [9] in Fig. 4. As is shown in Fig. 4 (a), the blank car in the red rectangle stands by the side of the road. Vista and EOT-WM w/o O.V. Traj. do not consider the other vehicle trajectories, and the blank car moves forward in the generated videos. In contrast, the blank car keeps still in the video generated by the proposed EOT-WM. In Fig. 4 (b), the truck is backing up on the road. However, Vista and EOT-WM w/o O.V. Traj. both generate the truck moving forward in the generated video. As for the proposed EOT-WM, the action of the truck is predicted correctly in the generated video. These cases demonstrate that the proposed EOT-WM can generate more realistic videos with ego and other vehicle trajectories.
@@
4.4 Ablation Study
To verify the effectiveness of the proposed modules, the ablation study for key components is conducted on the validation set of nuScenes dataset, and the results are given in Tab. 2. When all components are not used, the model is close to action-free only with initial frames as the context. As a result, the model obtains poor performance as 61.4 FID and 372.8 FVD. As for EOT-WM w/o VTR model, the performance is still poor, since the model cannot learn trajectories in BEV space well and other modules designed for VTR cannot work. For EOT-WM w/o AMGG model, we use C3D to extract trajectory video features, and the performance declines compared with the proposed EOT-WM due to the dissimilarity of the feature space. The ego vehicle trajectory is the most important condition for video generation. Therefore, EOT-WM w/o E. V. Traj. model cannot generate realistic videos. With other vehicle trajectories to control the other vehicles in the scene, the proposed EOT-WM model can achieve better FID and FVD compared with EOT-WM w/o O. V. Traj. model. All above results demonstrate the effectiveness of the proposed modules.
4.5 Effect of difference text types
The original CogvideoX [43] is trained to use the text to control video generation, and other advanced video generation models [11, 23] are also text-to-video models. When building the world model based on such text-to-video models, it is unadvisable to remove the text element since the capability of the original text-to-video model will be damaged. As can be observed in Tab. 3, while the proposed EOT-WM does not use the text, FID and FVD achieved decline obviously. Then, we adopt three types of texts to control the video generation model. The performances obtained by these types of texts are similar, demonstrating that the video generation model can be trained to adapt different types of texts. Among them, EOT-WM using the action description achieves the best performance since the caption contains action information. However, considering the capability of generating novel scenes, we choose the scene description for EOT-WM as the action description will conflict with the novel trajectories.
4.6 Generating novel scenes
To verify the capability of the proposed EOT-WM to generate novel scenes, several instances generated with novel trajectories are given in Fig. 6. As can be observed in Fig. 6 (a), the ego vehicle turns right and the bus moves forward in the video generated with groundtruth ego and other vehicle trajectories in the dataset. While we use the self-produced ego vehicle trajectory that represents turning left in Fig. 6 (b), the ego vehicle is controlled to turn left in the video generated. Moreover, we give the novel other vehicle trajectory meaning stop in Fig. 6 (c), the bus in the generated video stands by. Above results demonstrate that the proposed EOT-WM can generate novel scenes according to novel trajectories.
4.7 Evaluating Trajectory Controllability
To evaluate trajectory controllability, we propose a metric based on control latent similarity in the video latent space. As is shown in Fig. 7, we use the conditional trajectory video to generate predicted trajectory video latents in TiDiT. The predicted trajectory video latents can be decoded into predicted trajectory video, which is similar to the conditional trajectory video. Moreover, given another irrelevant trajectory video, the metric calculated between the predicted trajectory video latents and the conditional trajectory video latents is obviously smaller than that calculated between the predicted trajectory video latents and the irrelevant trajectory video latents. These results demonstrate the effectiveness of the proposed latent-based metric.
@@ 4.7 Evaluating Trajectory Controllability
Table 3: Effect of difference texts used in EOT-WM.
@@ 5 Conclusion
In the future work, we will explore the manner to control the other vehicles more precisely and the quantitative metric to evaluate the consistency between the given trajectory and the generated video.